OpenStudy (anonymous):

A pyramid has a height of 5 in. and a surface area of 90 in2. Find the surface area of a similar pyramid with a height of 10 in. Round to the nearest tenth, if necessary. (1 point) 360 in2 180 in2 22.5 in2 3.6 in2

OpenStudy (anonymous):

This is just a proportion problem: \[\frac{ 5 }{ 90 } = \frac{ 10 }{ x } \rightarrow 5x = 900 \rightarrow x=\frac{ 900 }{ 5 }\]

OpenStudy (anonymous):

So SA = 180 in^2 for a height of 10.

OpenStudy (anonymous):

thanks. do you know this A rectangular prism has a width of 92 ft and a volume of 240 ft2. Find the volume of a similar prism with a width of 23 ft. Round to the nearest tenth, if necessary. (1 point) 3.8 ft3 60 ft3 15 ft3 10.4 ft3

OpenStudy (anonymous):

So basically it's another proportion: \[\frac{ 92 }{ 240 } = \frac{ 23 }{ X }\] Solve for x to get the ans.

OpenStudy (anonymous):

D( x ) x = 0 x = 0 x = 0 x in (-oo:0) U (0:+oo) 92/240 = 23/x // - 23/x 92/240-(23/x) = 0 92/240-23*x^-1 = 0 -23*x^-1 = -23/60 // : -23 x^-1 = 1/60 -1 < 0 1/(x^1) = 1/60 // * x^1 1 = 1/60*x^1 // : 1/60 60 = x^1 x = 60 x = 60